# Complex Physics-Informed Neural Network

### v1

@@ 4.2 Diffusion-reaction Equation
where
$q(x,y)=(k^{2}-2(a\pi)^{2})\sin(a\pi x)\sin(a\pi y).$ (32)
We set $k=1$ and $\Omega=[-1,1]\times[-1,1]$ and the exact solution to the equation is
$u(x,y)=\sin(a\pi x)\sin(a\pi y).$ (33)
In this study, we focus on two different settings for parameter $a\in\{3,6\}$ in the forcing term $q(x,y)$, where the corresponding exact solution can be viewed in Fig. 5. For both cases, we compare the compleX-PINN with the traditional PINN and RBA.
All PINN-based models employ a 5-layer network architecture, each layer consisting of 60 neurons. The compleX-PINN uses a Cauchy layer with 1000 neurons. Training is conducted using the Adam optimizer for 10k iterations with a learning rate of $1\times 10^{-4}$. Furthermore, we set $N_{f}=5000$ for each model. Boundary conditions are implemented using a hard constraint formulated as $\hat{u}=(x^{2}-1)(y^{2}-1)\hat{u}_{NN}$ [46, 47, 42]. As a result, only the residual loss (5) is enforced, and no boundary training points are included.
Figure 5: The exact solutions for the 2D Helmholtz equation when $a=3$ (left) and $a=6$ (right).
The results for each model can be visualized in Fig. 6. For $a=3$, both PINN and RBA exhibit significant errors spread across the domain, while compleX-PINN accurately predicts the solution across nearly the entire domain, with only small localized errors. Moreover, the maximum error for compleX-PINN ($2.11\times 10^{-3}$) is substantially lower than those of PINN ($9.02\times 10^{-2}$) and RBA ($3.01\times 10^{-2}$).
Moreover, for the $a=6$ case, point-wise error comparisons across PINN, RBA, and compleX-PINN highlight the better accuracy of compleX-PINN. The PINN prediction exhibits regions of significant error, particularly near the boundaries and in specific interior areas, with a maximum error of $0.383$. The RBA method reduces these errors but still shows localized high-error regions, achieving a maximum error of $0.146$. In contrast, compleX-PINN demonstrates substantially improved performance with a significantly lower maximum error of $1.97\times 10^{-2}$.
The relative $L^{2}$ error history is shown in Fig. 7. We can see the relative $L^{2}$ errors for compleX-PINN decrease faster to lower values, compared with PINN and RBA. These results underscore the effectiveness of compleX-PINN in achieving higher accuracy and better error control compared to the baseline methods.
PINN ($a=3$): $9.02\times 10^{-2}(L^{\infty})$
PINN ($a=6$): $3.83\times 10^{-1}(L^{\infty})$
RBA ($a=3$): $3.01\times 10^{-2}(L^{\infty})$
RBA ($a=6$): $1.46\times 10^{-1}(L^{\infty})$
compleX-PINN ($a=3$): $\bm{2.11\times 10^{-3}}(L^{\infty})$
compleX-PINN ($a=6$): $\bm{1.97\times 10^{-2}}(L^{\infty})$
Figure 6: Point-wise errors for traditional PINN (first row), RBA (second row), and compleX-PINN (third row), respectively. The first column stands for the results for $a=3$ and the second is for $a=6$. compleX-PINN receives a lower $L^{\infty}$ error for both cases, compared with PINN and RBA.
Figure 7: Relative $L^{2}$ error history of PINN, RBA, and compleX-PINN for the 2D Helmholtz equation for $a=3$ and $6$. CompleX-PINN converges faster to a lower value. E.g., when $a=3$, the $L^{2}$ error is reduced to below $0.01$ for about 20k iterations, while the errors for the other two models can only archive the lowest value around $0.03$.
4.4 Reaction equation
The reaction problem is a hyperbolic PDE used to model chemical reactions, formulated with periodic boundary conditions as follows:
$u_{t}-\rho u(1-u)=0,\quad x\in[0,2\pi],\ t\in[0,1],$ (34)
$u(0,t)=u(2\pi,t),\quad\ t\in[0,1],$ (35)
$u(x,0)=u_{0}(x),\quad x\in[0,2\pi].$ (36)
The system has an analytical solution in the form:
$u(x,t)=\frac{u_{0}(x)e^{\rho t}}{u_{0}(x)e^{\rho t}+1-u_{0}(x)}.$ (37)
We define the initial condition as:
$u_{0}(x)=\exp\left(-\frac{(x-\pi)^{2}}{2(\pi/4)^{2}}\right).$ (38)
We evaluate the performance of PINN, RBA, compleX-PINN, and compleX-RBA for $\rho\in\{5,10,15,20\}$. We set $N_{f}=6000$ and $N_{0}=N_{b}=150$. PINN is set to be a 5-layer network with 80 neurons in each hidden layer and compleX-PINN has 1200 neurons for the Cauchy layer. Training is conducted using the Adam optimizer with a learning rate of $5\times 10^{-4}$ for 10k iterations. The relative $L^{2}$ error for all models are shown in Fig. 8 and Table 2. This figure illustrates that solving the PDE becomes increasingly challenging as the parameter $\rho$ grows larger. Notably, compleX-PINN consistently outperforms PINN and RBA, demonstrating its ability to handle higher stiffness in the reaction equation. Besides, compleX-RBA can reduce the relative $L^{2}$ errors to a lower value. This further demonstrates the potential of our model, suggesting that its performance can be enhanced by incorporating techniques originally proposed for PINNs, consistent with the conclusion in Section 4.1.
Figure 8: Relative $L^{2}$ error history of PINN, RBA, compleX-PINN, and compleX-RBA for the reaction equation with $\rho\in\{5,10,15,20\}$. In detail, PINN fails to train for large $\rho\in\{10,15,20\}$. Although RBA can make the relative $L^{2}$ errors to be below $1\times 10^{-1}$, compleX-PINN consistently outperforms them. The highest relative $L^{2}$ of compleX-PINN at 10k iteration for different $\rho$ values is about $1.52\times 10^{-2}$ and the lowest can archieve $7.04\times 10^{-3}$. Also, when we use the RBA technique to compleX-PINN, compleX-RBA can reduce the errors to lower values.
Table 2: Relative $L^{2}$ errors of the reaction equation for different models with $\rho\in\{5,10,15,20\}$ after 10k iterations.
$\rho$ value PINN RBA compleX-PINN compleX-RBA
5 $6.90\times 10^{-2}$ $3.31\times 10^{-2}$ $7.04\times 10^{-3}$ $4.23\times 10^{-3}$
10 $9.64\times 10^{-1}$ $3.40\times 10^{-2}$ $1.38\times 10^{-2}$ $4.75\times 10^{-3}$
15 $9.65\times 10^{-1}$ $4.15\times 10^{-2}$ $1.52\times 10^{-2}$ $5.07\times 10^{-3}$
20 $9.91\times 10^{-1}$ $6.42\times 10^{-2}$ $1.18\times 10^{-2}$ $6.23\times 10^{-3}$
4.5 Convection equation
The convection equation [48, 36, 28, 49] describes the transport of a conserved scalar field by a velocity field over time. This equation is widely applied in fluid dynamics and other related fields. We consider the following equation:
$u_{t}+\beta u_{x}=0,\quad(x,t)\in[0,2\pi]\times[0,1],$ (39)
$u(0,t)=u(2\pi,t),\quad t\in[0,1],$ (40)
$u(x,0)=\sin(x),\quad x\in[0,2\pi],$ (41)
with the corresponding exact solution:
$u(x,t)=\sin(x-\beta t),$ (42)
which oscillates periodically between $-1$ and $1$ and $\beta$ denotes the constant velocity parameter.
Previous studies have successfully trained PINNs for $\beta=50$, including approaches such as importance sampling [28], transformer-based PINN [36], and sequence-to-sequence learning [48]. A more recent work [49] extended this to $\beta=100$, which improves the accuracy of PINNs for time-dependent problems by exactly enforcing temporal continuity between sequential time segments through a solution ansatz. The lowest relative $L^{2}$ error recorded in Ref. [49] for $\beta=50$ and $100$ is $3.75\times 10^{-3}$ and $6.28\times 10^{-3}$, respectively.
However, as $\beta$ increases further, the complexity of the problem increases significantly, making it increasingly difficult for traditional PINNs to maintain accurate predictions. The aforementioned works do not extend to higher $\beta>100$. To address this, we extend the research by testing compleX-PINN for higher values of $\beta$, specifically for $\beta\in\{50,80,90,100,120,150\}$.
We use 2500 neurons for the Cauchy layer and set $N_{f}=50\times\beta$ and $N_{0}=N_{b}=100$. The compleX-PINN is trained using the Adam optimizer with a learning rate of $5\times 10^{-4}$ over 100k iterations.
The relative $L^{2}$ error history is shown in Fig. 9, and the final performance after training is summarized in Table 3. The performance for $\beta=50$ and $\beta=100$ exceeds that reported in Ref. [49]. Furthermore, we extend the analysis to $\beta=120$ and $\beta=150$, both of which achieve relative $L^{2}$ errors below $1\times 10^{-2}$.
@@ 4.5 Convection equation
Figure 9: Relative $L^{2}$ error history of compleX-PINN for the convection equation with $\beta\in\{50,80,90,100,120,150\}$.
Table 3: Final relative $L^{2}$ error of compleX-PINN for the convection equation after 100k iterations with different $\beta$.
$\beta=50$ $\beta=80$ $\beta=90$ $\beta=100$ $\beta=120$ $\beta=150$
Relative $L^{2}$ error $1.69\times 10^{-3}$ $3.69\times 10^{-3}$ $3.59\times 10^{-3}$ $4.82\times 10^{-3}$ $7.54\times 10^{-3}$ $7.66\times 10^{-3}$
5 Conclusion
In this paper, we introduce compleX-PINN, which is a single-layer network that utilizes the Cauchy integral formula. Complementing our theoretical advances, the empirical results show that compleX-PINN can achieve a much lower relative $L^{2}$ error with fewer iterations. Also, the number of parameters in compleX-PINN is much smaller than in the traditional PINN. This enhanced efficiency and accuracy validate the efficacy of our theoretical contributions and highlight the practical advantages of compleX-PINN.

### v2

@@ 4 Numerical Experiments
where the source term $q(x,y)$ is defined as:
$q(x,y)=k^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)-(a_{1}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)-(a_{2}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y).$ (20)
We set $k=1$ and $\Omega=[-1,1]\times[-1,1]$. The exact solution to the equation is given by
$u(x,y)=\sin(a_{1}\pi x)\sin(a_{2}\pi y),$ (21)
where the parameters $a_{1}$ and $a_{2}$ control the spatial frequency components of the solution.
4.1.1 Comparison with baseline models
We evaluate compleX-PINN against several baseline models: RBA-PINN [46], PIKAN [37, 38](with $a_{1}=1$ and $a_{2}=4$), and BsPINN [41] (with $a_{1}=a_{2}=8$). Each model is tested under its original PDE configuration, and we strictly follow the training protocols specified in the respective studies. This approach ensures fair, architecture-focused comparisons by eliminating confounding factors such as differences in hyperparameter tuning.
RBA-PINN Comparison (Case 4.1.1a):
We construct a compleX-PINN model with 500 neurons in the Cauchy layer. For comparison, the RBA-PINN employs a 5-layer fully connected neural network with 128 neurons per layer. To enforce boundary conditions, we use a hard constraint formulation:
$\hat{u}=(1-x^{2})(1-y^{2})\hat{u}_{\mathcal{NN}}$
where $\hat{u}_{\mathcal{NN}}$ is the raw network output and $\hat{u}$ is the final prediction satisfying the PDE boundary conditions.
To ensure comparability, we adopt the same number of training points $N_{f}=256,000$ and the same number of iterations as in RBA-PINN [46]. Specifically, we train using the Adam optimizer for 20,000 iterations with a learning rate of 0.005, followed by 1,000 iterations of L-BFGS for fine-tuning. An exponential learning rate scheduler with decay rate 0.7 is applied every 1,000 steps. All training details—including learning rates, scheduling, and network structure—closely follow the publicly available code provided by RBA-PINN.
Given that RBA-PINN has demonstrated state-of-the-art performance on this benchmark, we compare directly with its reported results. The relative $L^{2}$ error history during Adam training is shown in Figure 1, where the blue curve represents the RBA-PINN output using its official code. After the Adam phase, compleX-PINN achieves a relative $L^{2}$ error of $2.39\times 10^{-5}$, significantly outperforming the RBA-PINN’s corresponding error of $1.18\times 10^{-4}$. After the L-BFGS fine-tuning phase, compleX-PINN further reduces the error to $7.51\times 10^{-7}$, nearly one order of magnitude lower than RBA-PINN’s reported post-L-BFGS error of $5.18\times 10^{-6}$. The final prediction results are visualized in Figure 2.
Figure 1: Relative $L^{2}$ error history of the 2D Helmholtz equation using the Adam optimizer (red: compleX-PINN; blue: RBA-PINN), followed by 1,000 iterations of the L-BFGS optimizer. A dashed vertical line indicates the transition from Adam to L-BFGS. The Adam optimizer is configured with a learning rate of 0.005 and an exponential decay scheduler (decay rate: 0.7; decay step: 1,000). Although the large initial learning rate causes some oscillation in the first 5,000 iterations, the relative $L^{2}$ error stabilizes as the learning rate decays. The subsequent L-BFGS phase further refines the solution.
Figure 2: Prediction and absolute point-wise errors of compleX-PINN for the 2D Helmholtz equation. The results are obtained after fine-tuning with the L-BFGS optimizer. The final relative $L^{2}$ error achieved is $7.51\times 10^{-7}$.
PIKAN Comparison (Case 4.1.1b):
To further validate the performance of compleX-PINN, we compare it with PIKAN [38] using the same training protocol reported in [38]. The PIKAN model consists of a single hidden layer with 10 neurons, the number of intervals $G=5$, and spline order of $p=3$. The architecture uses a multi-grid training strategy where the number of grid intervals doubles every 600 iterations. This results in a total of 750 parameters as reported in Ref. [37]. Accordingly, we configure compleX-PINN with 110 neurons in the Cauchy layer, resulting in a total of 771 parameters, comparable to PIKAN.
We use the same number of training points ($N_{f}=2401$, $N_{b}=200$), the same loss weights ($\lambda_{F}=0.01$, $\lambda_{B}=1$), and adopt the L-BFGS optimizer as done in Ref. [37]. All models are trained for exactly 1,800 iterations. With identical training data and hyperparameters, compleX-PINN achieves a relative $L^{2}$ error of $6.58\times 10^{-2}$, substantially outperforming the reported PIKAN result of $4.76\times 10^{-1}$ by nearly an order of magnitude.
BsPINN Comparison (Case 4.1.1c):
We now evaluate compleX-PINN against BsPINN on a more challenging 2D Helmholtz problem with parameters $a_{1}=a_{2}=8$, consistent with the setup in Ref. [41]. The BsPINN model uses a 256-16 binary configuration with five hidden layers organized hierarchically. Specifically, the first hidden layer contains one block of 256 neurons; the second layer has two blocks of 128 neurons each; the third has four blocks of 64 neurons; the fourth has eight blocks of 32 neurons; and the fifth has sixteen blocks of 16 neurons. This architecture results in 124,929 total parameters. In contrast, our compleX-PINN model reuses the architecture from Case 4.1.1 (500 neurons in the Cauchy layer), which has only 3,501 parameters—approximately 3% of BsPINN’s size—highlighting the efficiency of our design.
For both models, we use $N_{f}=6561$ interior training points, $N_{b}=320$ boundary points, and set loss weights as $\lambda_{F}=1$, $\lambda_{B}=100$. The models are trained using the Adam optimizer for 40,000 iterations with a learning rate of $1\times 10^{-3}$. Under this setup, compleX-PINN achieves a relative $L^{2}$ error of $5.35\times 10^{-5}$, significantly outperforming BsPINN’s best reported result of $1.07\times 10^{-3}$. The relative error history and heatmap of pointwise errors for compleX-PINN are shown in Figures 3 and 4, respectively.
@@ 4.1.1 Comparison with baseline models
Figure 3: Relative $L^{2}$ error and $L^{\infty}$ norm history of compleX-PINN for the 2D Helmholtz equation ($a_{1}=a_{2}=8$).
Figure 4: Prediction and absolute point-wise errors of compleX-PINN for the 2D Helmholtz equation when $a_{1}=a_{2}=8$. The final relative $L^{2}$ error after training is $5.35\times 10^{-5}$.
4.1.2 Robustness of the trainable parameters in Cauchy layers
Recall that the Cauchy activation function introduced in (8) contains trainable parameters ${\mu_{1},\mu_{2},d}$, which are manually initialized. By default, we initialize each of these parameters to 0.1. However, it is important to investigate whether this initialization impacts model performance.
In this section, we examine the robustness of these parameters by varying their initial values among $\{0.01,0.5,1.0\}$. Testing all possible combinations of $\{\mu_{1},\mu_{2},d\}$ would be computationally intensive. To balance thoroughness with feasibility, we adopt a controlled perturbation strategy: we fix two parameters at their default value (0.1) and vary the third. This allows us to isolate the effect of each individual parameter’s initialization on the model’s performance.
All experiments follow the same training configuration described in Case 4.1.1a, and we restrict our analysis to the Adam training phase.
Table 1: Robustness of model performance with respect to the initialization of Cauchy activation function parameters ${\mu_{1},\mu_{2},d}$. Each parameter is varied individually while the other two are fixed at 0.1. Reported values are the relative $L^{2}$ errors after 20,000 Adam training iterations (mean ± standard deviation over 5 trials). "NaN" indicates numerical instability during training.
Parameter Initialization Value
0.01 0.5 1.0
$\mu_{1}$ (fixed: $\mu_{2}=0.1,d=0.1$) $(2.38\pm 0.18)\times 10^{-5}$ $(2.39\pm 0.47)\times 10^{-5}$ $(2.58\pm 0.41)\times 10^{-5}$
$\mu_{2}$ (fixed: $\mu_{1}=0.1,d=0.1$) $(2.69\pm 0.22)\times 10^{-5}$ $(2.99\pm 0.57)\times 10^{-5}$ $(3.47\pm 0.83)\times 10^{-5}$
$d$ (fixed: $\mu_{1}=0.1,\mu_{2}=0.1$) NaN $(3.38\pm 0.46)\times 10^{-5}$ $(7.85\pm 0.16)\times 10^{-5}$
As shown in Table 1, variations in the initial values of $\mu_{1}$ and $\mu_{2}$ result in relatively consistent model performance, with only minor fluctuations in relative $L^{2}$ error. This suggests that the model is robust to the initialization of these two parameters. In contrast, the initialization of $d$ has a more pronounced impact. Specifically, setting $d=0.01$ leads to numerical instability (NaN results), while larger values increase the relative error. This sensitivity is expected, as $d$ controls the "width" of the Cauchy activation function. As a scale parameter, small values of $d$ sharply localize the activation, potentially leading to gradient instability and poor generalization. These findings highlight the importance of avoiding near-zero initializations for $d$ to ensure stable and effective training.
4.2 3D Helmholtz equation
We now extend our investigation to the 3D Helmholtz equation. Specifically, we consider the following problem:
$\Delta u+k^{2}u=q(x,y,z),\quad(x,y,z)\in\Omega,$
$u(x,y,z)=0,\quad(x,y,z)\in\partial\Omega,$
where $k=1$, and the analytical solution is given by
$u(x,y,z)=\sin(2\pi x)\sin(2\pi y)\sin(2\pi z),$
with the source term $q(x,y,z)$ inferred accordingly.
All models incorporate a hard constraint for the boundary condition, expressed as $\hat{u}=(1-x^{2})(1-y^{2})(1-z^{2})\hat{u}_{\mathcal{NN}}$, and use $N_{f}=10{,}000$ residual training points. Training is performed using the Adam optimizer for 20k iterations with the same learning rate and exponential decay scheduler as described in Case 4.1.1a.
The compleX-PINN employs 500 neurons in the Cauchy layer. For baseline models, hyperparameter tuning is conducted as described at the beginning of Section 4, and the best-performing configurations are selected accordingly: RBA-PINN uses $D_{\text{RBA}}=4$ and $N_{\text{RBA}}=50$; PIKAN uses $G=4$ and $N_{\text{KAN}}=5$; BsPINN adopts a 5-layer hierarchical architecture with block sizes halving from 256 to 16, resulting in a 256–16 configuration.
Table 2 summarizes the number of trainable parameters and the GPU time per 100 training iterations for each model. Notably, compleX-PINN and PIKAN have significantly fewer parameters than RBA-PINN and BsPINN, highlighting their advantage in model compactness and memory efficiency. While compleX-PINN has more parameters than PIKAN, it requires substantially less training time per 100 iterations, demonstrating superior computational efficiency.
Table 2: Number of trainable parameters and GPU training time per 100 iterations for different models. This comparison highlights the computational and memory efficiency of compleX-PINN relative to other baseline methods.
Model Number of Parameters GPU Time (s/100 iters)
compleX-PINN 4,001 5.23
RBA-PINN 7,901 3.05
BsPINN 125,185 4.81
PIKAN 405 11.70
Table 3 reports the numerical performance under two evaluation criteria: (i) the same number of training iterations (20k), and (ii) equivalent total training time. After 20k iterations, compleX-PINN achieves a relative $L^{2}$ error of $5.54\times 10^{-4}$. PIKAN has the second smallest $L^{2}$ error of $7.27\times 10^{-4}$ after 20k iterations, despite its longer training time. When training time is equalized, BsPINN and RBA-PINN attain relative $L^{2}$ errors of $9.94\times 10^{-4}$ and $1.03\times 10^{-3}$, respectively. Overall, compleX-PINN consistently outperforms other methods and yields the best performance across both evaluation metrics, even when trained under time constraints.
Table 3: Relative $L^{2}$ errors and $L^{\infty}$ norms for each model under two evaluation settings—fixed number of iterations (20k) and equivalent computational time.
Model (total 20k iter.) Relative $L^{2}$ Norm $L^{\infty}$ Norm
Same Iter. Same Comput. Time Same Iter. Same Comput. Time
compleX-PINN $5.54\times 10^{-4}$ $9.27\times 10^{-4}$
RBA-PINN $2.71\times 10^{-3}$ $1.03\times 10^{-3}$ $3.13\times 10^{-2}$ $3.01\times 10^{-2}$
BsPINN $1.03\times 10^{-3}$ $9.94\times 10^{-4}$ $4.25\times 10^{-3}$ $4.18\times 10^{-3}$
PIKAN $7.27\times 10^{-4}$ $3.58\times 10^{-3}$ $2.72\times 10^{-3}$ $5.07\times 10^{-3}$
4.3 3D heat equation
We now evaluate model performance on the 3D heat equation given by:
$\Delta u(\mathbf{x},t)=12\pi^{2}u_{t},\quad\mathbf{x}\in\Omega:=[0,1]^{3},t\in T:=[0,10],$ (22)
$u(\mathbf{x},t)=0,\quad\mathbf{x}\in\partial\Omega,t\in T,$ (23)
$u(\mathbf{x},0)=\sin(2\pi x_{1})\sin(2\pi x_{2})\sin(2\pi x_{3}),\quad\mathbf{x}\in\Omega.$ (24)
The exact solution is:
$u(\mathbf{x},t)=e^{-t}\sin(2\pi x_{1})\sin(2\pi x_{2})\sin(2\pi x_{3}).$ (25)
To enforce initial and boundary conditions, we apply a hard constraint of the form:
$\hat{u}=t\prod_{i=1}^{3}x_{i}(1-x_{i})\cdot\hat{u}_{\mathcal{NN}}+\prod_{i=1}^{3}\sin(2\pi x_{i}).$
We use $N_{f}=10,000$ residual training points across space and time. The compleX-PINN is configured with 300 neurons in its Cauchy layer. For comparison, the baseline models are optimized with the following architectures based on prior tuning: RBA-PINN uses 4 hidden layers with 80 neurons each; BsPINN adopts a 256-16 block-wise configuration; and PIKAN uses $N_{\text{KAN}}=10$ and $G=2$.
All models are trained using the Adam optimizer with an initial learning rate of $5\times 10^{-3}$, decayed exponentially by a factor of 0.85 every 1,000 iterations, until below $1\times 10^{-5}$. Figure 5 presents the training history of the relative $L^{2}$ error, and Table 4 summarizes each model’s performance under two evaluation protocols: fixed number of training iterations (20,000) and fixed total training time (matched to compleX-PINN).
Figure 5: Relative $L^{2}$ error history of compleX-PINN (red), RBA-PINN (blue), BsPINN (cyan), and PIKAN (magenta) on the 3D heat equation. The left plot shows error evolution over 20,000 training iterations. The right plot presents the error history when models are trained for the same total GPU time as compleX-PINN.
Table 4: 3D heat equation: Relative $L^{2}$ errors and $L^{\infty}$ norms for each model under two evaluation settings—fixed number of iterations (20k) and equivalent computational time.
Model (total 20k iter.) Relative $L^{2}$ Norm $L^{\infty}$ Norm GPU Time (per 100 iter.)
Same Iter. Same Comput. Time Same Iter. Same Comput. Time
compleX-PINN $2.73\times 10^{-4}$ $7.67\times 10^{-4}$ 4.25s
RBA-PINN $1.16\times 10^{-3}$ $1.06\times 10^{-3}$ $2.27\times 10^{-3}$ $2.23\times 10^{-3}$ 3.03s
BsPINN $1.75\times 10^{-2}$ $1.77\times 10^{-2}$ $3.58\times 10^{-2}$ $3.76\times 10^{-2}$ 5.05s
PIKAN $1.10\times 10^{-3}$ $3.12\times 10^{-3}$ $1.39\times 10^{-3}$ $4.19\times 10^{-3}$ 20.8s
Under identical iteration count, compleX-PINN achieves a relative $L^{2}$ error of $2.73\times 10^{-4}$, significantly outperforming the other three methods. This advantage persists even under equal runtime. For example, although RBA-PINN completes over 28,000 iterations within the same time (due to its faster per-iteration speed of 3.03s per 100 iterations), its final error remains at $1.06\times 10^{-3}$, still nearly 4× higher than compleX-PINN. PIKAN, despite achieving reasonable accuracy ($1.10\times 10^{-3}$) with 20k iterations, suffers from high computational cost (20.8s per 100 iterations), allowing only  4,100 iterations under equal runtime. Consequently, its error increases to $3.12\times 10^{-3}$, over 11× higher than compleX-PINN.
In terms of maximum point-wise error, compleX-PINN again demonstrates superiority, achieving the lowest $L^{\infty}$ norm of $7.67\times 10^{-4}$, underscoring its effectiveness in reducing localized spatiotemporal errors. These results highlight the strength of the Cauchy-layer-based architecture in enhancing both global and worst-case accuracy.
4.4 High-dimensional Poisson equation
We evaluate the performance of compleX-PINN on high-dimensional problems using the Poisson equation in five and ten dimensions. The governing equation is defined on the hypercube domain $\Omega:=[-1,1]^{d}$, with $d\in{5,10}$:
$-\Delta u$ $=f\quad\text{in }\Omega,$ (26)
$u$ $=g\quad\text{on }\partial\Omega,$ (27)
where the exact solution is given by $u(\mathbf{x})=\sum_{i=1}^{d}\sin(\pi x_{i})$ and the corresponding source term is $f(\mathbf{x})=\pi^{2}\sum_{i=1}^{d}\sin(\pi x_{i})$. This setup provides a controlled and scalable benchmark for assessing the accuracy and efficiency of the models.
For both the 5D and 10D cases, all models are trained under identical settings: $N_{f}=10,000$ interior collocation points, $N_{b}=500$ boundary points, and 50,000 training iterations using the Adam optimizer with an initial learning rate of $1\times 10^{-2}$. A learning rate scheduler with decay factor 0.85 and decay step 1,000 is applied, and the loss weights are fixed at $\lambda_{F}=1$ and $\lambda_{B}=100$. The compleX-PINN employs 200 neurons in the Cauchy layer. Baseline models are configured with their optimal architectures: RBA-PINN uses 3 hidden layers of 200 neurons; PIKAN adopts $N_{\text{KAN}}=15$ with number of intervals $G=9$; and BsPINN follows a 128-8 architecture. The results are summarized in Table 5 and visualized in Figure 6.
In the 5D case, compleX-PINN achieves a relative $L^{2}$ error of $1.66\times 10^{-5}$ and an $L^{\infty}$ norm of $4.09\times 10^{-3}$, significantly outperforming all baselines by one to two orders of magnitude. When extended to 10D, compleX-PINN maintains remarkable robustness, with its $L^{2}$ error increasing only slightly to $3.44\times 10^{-5}$ (a 2.1× growth) and its $L^{\infty}$ norm rising to $4.54\times 10^{-3}$ (1.1× growth). In contrast, baseline models experience much greater degradation from 5D to 10D. For instance, RBA-PINN’s $L^{\infty}$ norm increases by 4.1× (from $1.58\times 10^{-2}$ to $6.45\times 10^{-2}$), BsPINN’s by 5.9× (from $1.09\times 10^{-2}$ to $6.46\times 10^{-2}$), and PIKAN’s nearly doubles (from $1.57\times 10^{-1}$ to $3.29\times 10^{-1}$). These results underscore compleX-PINN’s ability to suppress error amplification in high-dimensional settings.
In terms of efficiency, compleX-PINN also demonstrates strong scalability. The GPU time per 100 iterations increases by only 33% from 4.05s (5D) to 5.38s (10D)—notably lower than the growth observed in RBA-PINN (45%), BsPINN (47%), and PIKAN (58%). These results demonstrate that compleX-PINN not only achieves high accuracy in high dimensions but also maintains strong computational efficiency and stability, effectively overcoming the key challenges of high-dimensional PDE solvers.
Table 5: 5D and 10D Poisson equations: Relative $L^{2}$ errors and $L^{\infty}$ norms for each model under two evaluation settings—fixed number of iterations (20k) and equivalent computational time.
Model (5D) (total 50k iter.) Relative $L^{2}$ Norm $L^{\infty}$ Norm GPU Time (per 100 iter)
Same Iter. Same Comput. Time Same Iter. Same Comput. Time
compleX-PINN $1.66\times 10^{-5}$ $4.09\times 10^{-3}$ 4.05s
RBA-PINN $2.77\times 10^{-4}$ $2.42\times 10^{-4}$ $1.58\times 10^{-2}$ $1.40\times 10^{-2}$ 2.86s
BsPINN $2.19\times 10^{-4}$ $2.49\times 10^{-4}$ $1.09\times 10^{-2}$ $1.14\times 10^{-2}$ 5.11s
PIKAN $2.64\times 10^{-3}$ $7.37\times 10^{-3}$ $1.57\times 10^{-1}$ $4.18\times 10^{-1}$ 21.2s
Model (10D) (total 50k iter.) Relative $L^{2}$ Norm $L^{\infty}$ Norm GPU Time (per 100 iter)
Same Iter. Same Comput. Time Same Iter. Same Comput. Time
compleX-PINN $3.44\times 10^{-5}$ $4.54\times 10^{-3}$ 5.38s
RBA-PINN $5.84\times 10^{-4}$ $5.29\times 10^{-4}$ $6.45\times 10^{-2}$ $6.83\times 10^{-2}$ 4.16s
BsPINN $4.19\times 10^{-4}$ $5.30\times 10^{-4}$ $6.46\times 10^{-2}$ $6.93\times 10^{-2}$ 7.52s
PIKAN $5.48\times 10^{-3}$ $1.67\times 10^{-2}$ $3.29\times 10^{-1}$ $8.48\times 10^{-1}$ 33.6s
Figure 6: Relative $L^{2}$ error history of compleX-PINN (red), RBA-PINN (blue), BsPINN (cyan), and PIKAN (magenta) for the Poisson equation in 5D (top row) and 10D (bottom row). The left column presents error trajectories over 50,000 training iterations, while the right column compares performance under equal GPU time budgets. compleX-PINN consistently demonstrates faster convergence and superior accuracy across both dimensional settings.
5 Conclusion and Future Work
We propose compleX-PINN, a novel single-layer physics-informed neural network architecture grounded in the Cauchy integral formula. Theoretically, this framework reduces parameter complexity while maintaining rigorous approximation guarantees, directly addressing the inefficiencies and depth-related limitations of conventional PINNs. Empirically, compleX-PINN achieves state-of-the-art accuracy and stability across a variety of high-dimensional PDE benchmarks, consistently delivering orders-of-magnitude lower relative $L^{2}$ and $L^{\infty}$ errors than leading alternatives such as RBA-PINN, BsPINN, and PIKAN. By uniting mathematical rigor with practical scalability, compleX-PINN offers a lightweight yet powerful paradigm for physics-informed machine learning.
Looking ahead, a key direction is the extension of compleX-PINN to non-smooth and discontinuous systems, such as PDEs with shock waves or sharp interfaces. While the current framework excels in smooth or analytic settings, its reliance on the analyticity assumptions inherent in the Cauchy integral formula may limit its direct applicability to such problems. This limitation indicates a potential future research direction. We will conduct deeper research in this area to explore effective solutions. Furthermore, deep or hierarchical variants of compleX-PINN—such as networks composed of multiple Cauchy layers—could enhance its expressiveness for multiscale phenomena without compromising training efficiency. Finally, integrating compleX-PINN with operator learning frameworks presents an exciting opportunity for emulating parametric systems efficiently, bridging the gap between classical numerical methods and modern data-driven approaches.